OpenStudy (anonymous):

3. use the distributive property to simplify the expression . ( 10 + 4y ) 1.2 ( 1point ) A. 5+2y b.20 + 8y c. 5+4y d. 10+2y

OpenStudy (ahsome):

Equation:\[1.2( 10 + 4y )\]Expand Equation:\[(1.2*10)+(1.2*4y)\]\[12+4.8y\]

OpenStudy (ahsome):

I think you got the equation wrong, I am not sure about the \(1.2\) but

OpenStudy (anonymous):

it's 1 ___ 2

OpenStudy (ahsome):

OH, like:\[\frac{1}{2}\]?

OpenStudy (anonymous):

yep

OpenStudy (ahsome):

Equation:\[\frac{1}{2}(10+4y)\]Expand Equation:\[(\frac{1}{2}∗10)+(\frac{1}{2}∗4y)\]\[5+2y\]